### v2

# IssueGuard: Real-Time Secret Leak Prevention Tool for GitHub Issue Reports

Abstract.
GitHub and GitLab are widely used collaborative platforms whose issue-tracking systems contain large volumes of unstructured text, including logs, code snippets, and configuration examples. This creates a significant risk of accidental secret exposure, such as API keys and credentials, yet these platforms provide no mechanism to warn users before submission. We present IssueGuard, a tool for real-time detection and prevention of secret leaks in issue reports. Implemented as a Chrome extension, IssueGuard analyzes text as users type and combines regex-based candidate extraction with a fine-tuned CodeBERT model for contextual classification. This approach effectively separates real secrets from false positives and achieves an F1-score of 92.70% on a benchmark dataset, outperforming traditional regex-based scanners. IssueGuard integrates directly into the web interface and continuously analyzes the issue editor, presenting clear visual warnings to help users avoid submitting sensitive data. The source code is publicly available at https://github.com/disa-lab/IssueGuard , and a demonstration video is available at https://youtu.be/kvbWA8rr9cU .
Software Security, Secret Detection, Developer Tools, GitHub, Language Model, Chrome Extension
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXccs: Security and privacy Software and application securityccs: Software and its engineering Software post-development issues
1. Introduction and Related Work
Developers frequently handle sensitive credentials like API keys, OAuth tokens, and private encryption keys. Accidental leakage of these secrets presents a growing security risk, especially in modern distributed software systems (Kreichman, 2021). The scale of this issue is significant; a 2022 report found over ten million new hardcoded secrets in public repositories (GitGuardian, 2024). While previous work has focused mainly on detecting secrets in source code and configuration files (Jones, 2023; Kreichman, 2021; Meli et al., 2019; Sinha et al., 2015), secrets also leak through unstructured channels such as issue reports, documentation, and developer discussions.
Existing research on secret detection has largely targeted source code (Meli et al., 2019; Sinha et al., 2015; Saha et al., 2020; Feng et al., 2022; Basak et al., 2023). Early approaches relying on regular expressions and entropy-based heuristics (Meli et al., 2019) often suffer from high false positive rates due to a lack of contextual understanding. While more recent machine learning-based approaches have moderately improved performance (Saha et al., 2020; Feng et al., 2022), an effective detection model does not necessarily translate into a practical, preventative solution. Current platforms like GitHub and GitLab offer secret scanning, but these typically occur after submission or within CI/CD pipelines (TruffleSecurity (2016); 7). This reactive approach leaves developers vulnerable to exposing secrets in public logs before remediation can occur. Furthermore, conventional scanners like TruffleHog often generate high rates of false positives, leading to developer alert fatigue. Crucially, no existing tool provides real-time feedback within the issue-writing interface itself across these major platforms.
To bridge this gap, we present IssueGuard, a browser-based tool that brings real-time secret detection directly into the issue editors of GitHub and GitLab. Building on the contextual classification framework by Ahmed et al. (Ahmed et al., 2025), IssueGuard employs a two-stage pipeline: extracting candidates via regex and classifying them using a fine-tuned CodeBERT model. This approach significantly reduces false positives (achieving 92.70% F1-score (Ahmed et al., 2025)) leveraging contextual understanding and provides immediate, in-browser warnings as developers type, preventing leaks before they are posted.
The main contributions of this work are:
  1. (1)
    We introduce IssueGuard, a Chrome extension that integrates a transformer-based secret detection model into GitHub and GitLab interfaces, enabling real-time, pre-submission leak prevention.
  2. (2)
    We conduct a comprehensive usability study with 50 participants, demonstrating the tool’s effectiveness and smooth integration into developer workflows.
The rest of the paper is organized as follows. Section 2 describes the system architecture. Section 3 assesses model effectiveness, real-time latency, and usability and Section 4 concludes the paper.
2. Methodology
In this section, we outline design and development of IssueGuard whose workflow is mentioned in Figure 1.
Figure 1. Workflow of IssueGuard
2.1. Core Detection Engine
IssueGuard is powered by the classification pipeline introduced in our prior work (Ahmed et al., 2025). This engine serves as the backend for our tool. We chose this pipeline as its underlying model, a fine-tuned CodeBERT, was shown to be effective at distinguishing real secrets from false positives in the environment of issue reports. The pipeline’s model was trained on a benchmark dataset introduced in our prior work (Ahmed et al., 2025). The dataset was constructed by collecting issue reports from GitHub repositories using keyword-based filtering (e.g., “key,” “token”), followed by applying 761 regular expressions to extract candidate secret strings. These candidates were annotated as either Secret or Non-sensitive (e.g., placeholders, redacted values, or dummy keys), resulting in 54,148 labeled instances, including 5,881 true secrets, split into 75% training, 10% validation, and 15% test sets. For semantic feature extraction, the engine employs the CodeBERT-base model (Feng et al., 2020), a bimodal pre-trained model for programming and natural languages chosen for its performance on mixed-language tasks and fast inference speed. A 200-character context window surrounding each candidate string is extracted from the issue report and concatenated with the candidate itself before being encoded by CodeBERT. The resulting embedding is passed to a connected classification head trained using cross-entropy loss to predict the binary Secret/Non-sensitive label. The fine-tuned model is then saved and integrated as the classification engine.
2.2. IssueGuard System Architecture and Workflow
IssueGuard follows a client–server architecture consisting of a Google Chrome extension as the client and a FastAPI-based backend service as the server. This design choice is driven by the goal to keep the browser extension responsive while performing expensive secret detection using a neural model. By offloading model inference to the backend, the extension remains fast and unobtrusive during issue writing. The tool is designed to integrate into developers’ existing workflows on GitHub and GitLab. Figure 2 illustrates the workflow. As developers write an issue report, IssueGuard performs background analysis and provides immediate feedback before the issue is submitted. The detection pipeline proceeds follows:
  1. (1)
    Text Monitoring: The Chrome extension monitors paste events inside the issue description editor on GitHub/GitLab.
  2. (2)
    Debounced Requests: To avoid communication during continuous typing, the extension applies a debouncing strategy.
  3. (3)
    Deferred Transmission: The current text is sent to the backend only after a short pause in user input.
  4. (4)
    Candidate Extraction: Upon receiving the text, the backend applies a set of 761 regular expressions to identify potential secret candidates.
  5. (5)
    Contextual Classification: Each candidate is paired with a 200-character context window and passed to the fine-tuned CodeBERT-based classifier.
  6. (6)
    False Positive Filtering: The classifier filters out benign strings such as placeholders, hashes, and redacted values.
  7. (7)
    Result Transmission: The backend returns confirmed secrets to the browser extension.
  8. (8)
    User Feedback: The extension highlights detected secrets in the editor and displays a tooltip warning, allowing users to correct the issue before submission.
Figure 2. Demonstration IssueGuard. As the user types, the extension sends text to the backend. Real secrets are highlighted in red, while regex-based positives such as placeholders are ignored.
Low latency is a requirement for interactive developer tools. To achieve this, IssueGuard incorporates several engineering optimizations. On the client side, debouncing reduces backend requests and limits network overhead. On the server side, FastAPI is used for its asynchronous and non-blocking execution model, allowing the backend to handle concurrent requests efficiently. Model inference is executed in a thread pool so that expensive computations do not block request handling. When running on GPU-enabled systems, the backend uses mixed-precision (FP16) inference (Micikevicius et al., 2018), which reduces memory usage and improves throughput without affecting prediction quality. In addition, a Least Recently Used (LRU) cache stores inference results. This allows repeated analyses of similar text to be returned, further improving responsiveness during iterative editing.
We evaluate the system’s performance by deploying the backend on a machine with an AMD Ryzen 3700G CPU, 16 GB RAM, and an NVIDIA RTX 3060 GPU. The architecture supports local execution within environments, ensuring that issues with sensitive information remain under the user’s control, avoiding the privacy concerns associated with cloud-hosted security.
3. Evaluation
In this section, we evaluate the effectiveness of IssueGuard. We assess three key aspects: 1) the classification effectiveness of its underlying model compared to state-of-the-art LLMs, 2) the real-time performance and latency of the tool itself, and 3) its practical usability and developer acceptance, measured via a user study.
3.1. Model Effectiveness
As outlined in the methodology, IssueGuard integrates the CodeBERT-based classification pipeline from our prior work (Ahmed et al., 2025). While newer Large Language Models (LLMs) such as StarCoder and specialized encoders like StarPII have shown promise in code tasks and personal information detection, our evaluation confirms that fine-tuned CodeBERT remains the optimal choice for real-time secret detection in issue reports.
To validate this choice, we compared our model against several other models, including StarCoder (15B) (Li et al., 2023), StarPII (Allal et al., 2023), and RoBERTa-base. As shown in Table 1, our fine-tuned CodeBERT model achieves an F1-score of 92.70%, outperforming both the much larger StarCoder (89.01%) and the PII-specialized StarPII (90.24%). While StarPII shows high precision, it suffers from lower recall in the noisy context of issue descriptions. CodeBERT provides the best balance of precision and recall while remaining lightweight enough for low-latency inference.
Table 1. Comparative analysis of detection models
Model Precision Recall F1-Score
IssueGuard (CodeBERT) 92.49% 92.91% 92.70%
RoBERTa-base 91.10% 89.40% 90.24%
StarPII (BigCode) 93.05% 85.50% 89.11%
StarCoder (15B) 88.50% 89.50% 89.00%
Regex-only 6.80% 100.0% 12.80%
Furthermore, the model’s generalization was validated on a set of 178 real-world GitHub repositories. As shown in Table 2, the model achieved a macro-average F1-score of 81.60%, demonstrating a strong ability to generalize beyond its training data.
Table 2. Model performance on 178 real-world repositories (from (Ahmed et al., 2025)).
Class Precision Recall F1-score
Secret 50.98% 86.67% 64.20%
Macro Avg. 75.35% 92.45% 81.60%
3.2. Comparison with Standard Security Tools
Existing secret detection tools such as TruffleHog and Gitleaks are industry standards for post-commit remediation, detecting secrets after they have entered source code or logs. However, using these repository-level scanners in real-time prevention workflows leads to alert fatigue. As shown in Table 3, TruffleHog and Gitleaks achieve high recall (over 93% and 90%), making them well suited for security audits, but their regex-based approaches yield low precision (around 50%), which would cause frequent false alerts in an interactive setting. In contrast, IssueGuard is designed for pre-submission use and operates directly on issue reports, an attack surface not covered by existing scanners. By leveraging semantic context from a fine-tuned CodeBERT model, it filters out common false positives such as placeholder values, achieving over 92’% precision while maintaining high recall (F1-score: 92.70%). Thus, while traditional scanners protect repositories post-commit, IssueGuard provides a high-precision, pre-commit defense that prevents secrets from leaving the developer’s local environment.
Table 3. Performance comparison of IssueGuard with other secret scanners.
Tool Precision Recall F1-Score
IssueGuard 92.49% 92.91% 92.70%
TruffleHog (v3.0) 55.39% 93.94% 69.69%
Gitleaks (v8.18) 49.10% 90.20% 63.60%
3.3. Tool Performance and Latency
Beyond model accuracy, a real-time tool must be fast. As detailed in our methodology, IssueGuard’s architecture is highly optimized with asynchronous processing and caching. Latency was measured as the time from completing a typing event in the GitHub issue editor to the browser’s visual highlighting of detected secrets. To ensure a rigorous evaluation, these measurements were conducted using the test split defined in our dataset, ensuring a representative mix of short and long issue descriptions.
The results show an average end-to-end prediction time of 198 milliseconds (0.198 seconds). The core model inference step is extremely efficient, accounting for only 10.1 milliseconds (0.0101s) of this time. This demonstrates that the tool’s architecture is well-suited for uninterrupted real-time application.
3.4. User Study on Tool Usability
To evaluate practical utility, we conducted a user study with 50 participants (25 developers, 12 testers, 6 team leads, and 7 managers). Participants were recruited via internal company mailing lists and university developer forums to ensure a diverse range of experience (0–10+ years).
Participants were tasked with submitting 10 issue reports on GitHub, five using the IssueGuard extension and five manually, with instructions to identify potential secrets (e.g., AWS keys, private tokens) within the reports. Following the task, we administered a survey based on Kitchenham and Pfleeger’s guidelines (Kitchenham and Pfleeger, 2008). The results, summarized in Table 4, were highly positive. In particular, participants reported significantly higher confidence when using IssueGuard compared to manual inspection, noting that the tool reduced the risk of overlooking subtle or partially redacted secrets.
Overall, 90% of participants rated their confidence in the tool at 4 or higher (on a 5-point scale), while feature satisfaction was similarly strong, with 80% of users reporting being “Very Satisfied” with ease of use. Qualitative feedback further highlighted the tool’s “native” feel, with users emphasizing that unobtrusive highlights reduced the cognitive load associated with manually searching for sensitive information. Feature satisfaction (Figure 3) was similarly exceptional. A vast majority reported being ‘Very Satisfied’ with the tool setup (75%), ease of use (80%), execution time (78%), and highlights (72%). Even tool documentation, which is typically a lower-rated aspect in software tools, was highly praised, with 93% of users expressing satisfaction. Qualitative feedback from the open-ended questions verified the quantitative data. Participants frequently noted that the presence of real-time feedback increased their confidence in submitting issues publicly, suggesting that IssueGuard not only detects leaks but also improves developers’ perceived security assurance.
Table 4. User Study: Confidence in IssueGuard’s Detection
Confidence Rating (1-5) # of Participants Percentage
5 (Very High Confidence) 40 80.0%
4 (High Confidence) 5 10.0%
3 (Moderate Confidence) 4 8.0%
2 (Low Confidence) 1 2.0%
1 (Very Low Confidence) 0 0.0%
Total 50 100.0%
Figure 3. Participants’ satisfaction with various aspects of the IssueGuard tool.
4. Conclusion and Future Scope
In this paper, we presented IssueGuard, a practical tool for preventing accidental secret leakage in GitHub issue reports. By integrating a high-performance CodeBERT-based contextual classifier into a low-latency Chrome extension, IssueGuard brings real-time secret detection directly into developers’ issue submission workflows. Our evaluation demonstrates that the tool achieves high detection accuracy while maintaining responsiveness suitable for real-time use. Overall, this work shows how contextual secret detection can be effectively operationalized as a preventative, developer-facing security tool. As future work, we plan to extend IssueGuard’s deployment to additional collaboration platforms and browsers.